OpenStudy (anonymous):

I need a bit of explanation please. No direct answers. A water sprinkler has a range of 5 meters as shown. The jet of water from the sprinkler sweeps out at an angle of 85° as the nozzle turns. What is the area watered by the sprinkler? Use the value π = 3.14, and round your answer to one decimal place.

OpenStudy (anonymous):

A. 15.6 square meters B. 17.8 square meters C. 18.5 square meters D. 19.2 square meters

OpenStudy (anonymous):

@iGreen Could you please lend me a hand here?

OpenStudy (anonymous):

@Nnesha Would you mind to quickly explain what I have to do to find my answer? Please.

OpenStudy (anonymous):

@iambatman @ganeshie8 @TuringTest Please??

ganeshie8 (ganeshie8):

whats the area of circle with radius 5 ?

OpenStudy (anonymous):

what I have done so far is find the circumference. which my guess is that 5 is the radius

OpenStudy (anonymous):

but I have to make it into a 0.5 when I do this. Look, 2pi(0.5)

OpenStudy (anonymous):

that equals 3.14

ganeshie8 (ganeshie8):

the question is about area, not the circumference

OpenStudy (anonymous):

but then when put it into the equation to find the length of the arc it doesn't work out so well. I put 85/360 x 3.14

OpenStudy (anonymous):

I don't know about finding that. Maybe in the subject is called something else but I don't really get it.

ganeshie8 (ganeshie8):

First find the area of circle of radius : 5

OpenStudy (anonymous):

0.741?

OpenStudy (anonymous):

I'm looking at my notes and this is what I got.

OpenStudy (anonymous):

It says measure of central angle/360 x pi

OpenStudy (anonymous):

the measure of the central angle is 85. so I did 85/360 x pi

ganeshie8 (ganeshie8):

what formula are you using for area of circle ?

OpenStudy (anonymous):

It's called radian measure

ganeshie8 (ganeshie8):

look up area of circle formula in ur notes

OpenStudy (anonymous):

I don't have that. I was guessing that it would be one of the formulas I had. I wrote everything I went through in the lesson and it wasn't included

OpenStudy (anonymous):

okay I found it on google its a=rpi^2

OpenStudy (anonymous):

no A=pi(r)^2

OpenStudy (anonymous):

i got 0.785

ganeshie8 (ganeshie8):

yes Area = pi*r^2 plugin r = 5, what do you get ?

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

for some reason in the lesson, it was required that I made the radius into a decimal. so if the radius was 5 it'll be 0.5

ganeshie8 (ganeshie8):

Yes, thats the area of entire region interior to circle

ganeshie8 (ganeshie8):

no need for all that

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So what is my next step

ganeshie8 (ganeshie8):

we want the area corresponding 85 degrees only, so multiply that area of full curcle, 78.5, by the ratio 85/360

OpenStudy (anonymous):

oh okay thank you so much!!

OpenStudy (anonymous):

18.53

ganeshie8 (ganeshie8):

Area watered by the sprinkler is given by \[\dfrac{85}{360}\times 78.5\] simplify

OpenStudy (anonymous):

Thank you for your help! I appreciate it a lot! huge help :)

ganeshie8 (ganeshie8):

`18.5` is correct they want you round the answer to just one decimal place

ganeshie8 (ganeshie8):

yw:)